News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 23 February 2001 10.30am EST

During last week the observing schedule was replanned to include the Gamma Ray Burster GRB010214 which was accepted as a Target Of Opportunity (TOO) on 22 Feb following its detection by the Beppo SAX mission. A very fast response was possible for this TOO due to good mission planning and a favorable combination of the timing of the trigger, pass times and the load configuration. The TOO was triggered at 6.30am EST and the maneuver to the target was made at 5pm EST.

Chandra continued through the Spring 2001 eclipse season with nominal power and thermal performance.

Real-time procedures were completed last week on 16 Feb to conduct an ACIS Squeegee test, on 17 Feb to uplink the medium Earth eclipse patch and also on 17 Feb to retract the LETG. All procedures were completed as planned.

The replanned schedule including the TOO and the observations for the coming week shown below. The portion of the Hubble Deep Field North observation that was impacted by the TOO is expected to be added to a planned observation of the same target scheduled for next week.

------------------------------------------
HDF-N (interrupted)     ACIS-I       Feb 21 
TOO GRB010214           ACIS-S       Feb 22
HDF-N (resumed)         ACIS-I
Radiation Belt                       Feb 24 
NGC 3783                ACIS-S/HETG
Radiation Belt                       Feb 26
NGC 3783                ACIS-S/HETG  Feb 27 
Radiation Belt                       Mar 1
HDF-N                   ACIS-I
Radiation Belt                       Mar 3
Coma                    HRC-I        Mar 4
NGC2516                 HRC-I        Mar 5

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports